A point charge q0 that has a charge of 0.800 C is at the origin. A second particle q that has a charge of 2.00 μC and a mass of 0.0800 g is placed at x 0.800 m. 1) What is the potential energy of this system of charges? (Express your answer to three significant figures.) 1.8e-3 mJ Submit You currently have 7 submissions for this question. Only 10 submission are allowed. You can make 3 more submissions for this question. Your submissions: 1.8e-3x Computed value: 1.8e-3 Submitted: Sunday, January 27 at 6:47 PM Feedback: It appears you have made a power of ten error. 2) If the particle with charge q is released from rest, what will its speed be when it reaches x - 2.00 m? (Express your answer to three significant figures.) m/s Submit You currently have 0 submissions for this question. Only 10 submission are allowed. You can make 10 more submissions for this question.
Answer #1

1) Potential energy of the system of charges, U = kq0q/r
U = 9*10^9*0.8*10^-6*2*10^-6/0.8 = 18 mJ
Answer = 18

2) Kinetic energy of q = 1/2mv^2 = change in potential energy

U1 = 18 mJ
U2 = 9*10^9*0.8*10^-6*2*10^-6/2.0 = 7.2 mJ

1/2 *0.08*10^-3 *v^2 = (18-7.2)*10^-3 J
v = 16.43 m/s